OpenStudy (anonymous):

fin the highest common factor of 56 and 98

OpenStudy (anonymous):

the HCF or GCF is 14

OpenStudy (anonymous):

4*14=56 and 7*14=98

OpenStudy (asnaseer):

the best way to do these types of problems is to split each number into a product of its primes, so:\[56=2^3*7\]\[98=2*7^2\]then take the common factors between the two, in this case its 2 and 7. so HCF is 2*7=14

OpenStudy (anonymous):

i am talking bout the HCF

OpenStudy (anonymous):

HCF is the same thing as GCF, its just another way to say it.

OpenStudy (asnaseer):

HCF = highest common factor GCF = greatest common factor both refer to the same thing

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

cn i ask more questions

OpenStudy (asnaseer):

fire away...

OpenStudy (anonymous):

a hotem has 56 guests.....35 of the guest are male work out 35 out of 56 as a percanrage please shoe me ut working out

OpenStudy (asnaseer):

remember that 100% would be 56/56 so 35/56 should be (35*100)/56 %

OpenStudy (anonymous):

whats the answer them i do not get it at all

OpenStudy (anonymous):

divide 35 by 56 then times that decimal by 100 you'll get your answer

OpenStudy (anonymous):

o kl

OpenStudy (asnaseer):

100% means "everything" so if ALL the guests were male, then you know 100% of the guests are male since there are 56 guests, that means the ratio of males would be 56/56, or 1 in your problem, only 35 of the guests are male, so the proportion of guests that are male is 35/56 to turn this into a percentage you simply multiply this proportion by 100

OpenStudy (anonymous):

what is 625 as a percantage

OpenStudy (anonymous):

you almost got but it should have been a decimal. 62.5% would be the exact answer

OpenStudy (anonymous):

56/56 = 100% 35/56= .625 .625*100= 62.5 =62.5% is you answer